OpenStudy (anonymous):

Prove the identity (sin^4)x - (cos^4)x = (sin^2)x - (cos^2)x

OpenStudy (anonymous):

\[\sin ^{4}x - \cos ^{4}x = \sin ^{2}x - \cos ^{2}x\]

OpenStudy (turingtest):

apply difference of squares to the left once, and look for an identity

OpenStudy (anonymous):

sin^4x-cos^4x=(sin^2x+cos^2x )(sin^2x-cos^2x ).....(2) (sin^2x+cos^2x )=1.................(1) hence from 1 and 2 1((sin^2x-cos^2x )

OpenStudy (turingtest):

why use a double-angle formula? shivams' two-step answer is fine